If sin theta=3/4 and theta is an acute angle, what is the value of (tan theta)(cos theta)
Answer by stanbon(75887)

If sin theta=3/4 and theta is an acute angle, what is the value of (tan theta)(cos theta)
----------------
tan*cos=(sin/cos)*cos = sin(theta)
So, (tan(theta))(cos(theta))= sin(theta)=3/4
============
Cheers,
stan H.